OpenStudy (kayders1997):

Derivative of xy^1/2=x-2y

OpenStudy (abmon98):

Use product rule to differentiate xy^1/2. (f(x)g(x))=f'(x)g(x)+f(x)g'(x)->General form of the product rule d/dx(xy^1/2)=y^1/2+x1/2y^-1/2*dy/dx-->Chain rule in order to differentiate (y)^1/2 d/dx(x-2y)=1-2dy/dx y^1/2+x/2y^1/2*dy/dx=1-2dy/dx y^1/2-1=dy/dx(-2+x/2y^1/2) (y^1/2-1)(-2+x/2y^1/2)=dy/dx

OpenStudy (caozeyuan):

ok, I thought its asking partial of x and partial of y

OpenStudy (kayders1997):

Thank you Abmon!

OpenStudy (dannyo19):

wait//

OpenStudy (caozeyuan):

@DannyO19 , what?

OpenStudy (dannyo19):

The answer is 2sqrt y - 2y all over x + 4sqrt y

OpenStudy (dannyo19):

You need to use implicit derivation

OpenStudy (kayders1997):

Hmmmmm?

OpenStudy (dannyo19):

Make sure your doing your chain rule correctly and your multiplying the outside function utilizing the product and power rule by the inside function with respect to x, implicitly deriving y.

OpenStudy (kayders1997):

Ill try that thank you